OpenStudy (anonymous):

Find a ⋅ b. a = 7i - 4j, b = 4i + 3j

OpenStudy (e.cociuba):

Is that supposed to be: Find a - b. a = 7i - 4j b = 4k + 3j

OpenStudy (anonymous):

not that i know of, i copied and pasted it right from my worksheet, so idk

OpenStudy (e.cociuba):

* b = 4i + 3j

OpenStudy (e.cociuba):

This is Pre-Calculus. Am I correct?

OpenStudy (anonymous):

oh yes, i thought you meant to put a k. but yes this is.

OpenStudy (e.cociuba):

These are just 2 dimensional vectors, of the form "i" and "j". So, you are going to have to use (xa-xb), (ya-yb).

OpenStudy (anonymous):

so would the answer be 16?

OpenStudy (e.cociuba):

So, tell me. What would go into (xa-xb), (ya-yb) from these equations: a = 7i - 4j b = 4i + 3j

OpenStudy (anonymous):

(7i-4i) (-4j-3j)

OpenStudy (e.cociuba):

Correct now solve them.

OpenStudy (anonymous):

3i, -7j

OpenStudy (e.cociuba):

Correct. Good job :)

OpenStudy (e.cociuba):

Or if you wanted to leave it in coordinates, it would be (3, -7).

OpenStudy (anonymous):

<28, -12> 16 -40 <11, -1>

OpenStudy (anonymous):

can you show me how you did it? or would that take too long?

OpenStudy (anonymous):

did you just multiply each of the vectors?

OpenStudy (e.cociuba):

Correct, You said you got 16 before?

OpenStudy (anonymous):

yes,

OpenStudy (e.cociuba):

Okay, actually I think you made a typo which messed me up. Is it a+b?

OpenStudy (e.cociuba):

Because if we solve it any other way we are always going to come out with <3, -7> which is not correct. If the question is a+b then the last choice <11, -1> would have been correct.

OpenStudy (anonymous):

hen i look at it in the other window, it looks like a minus sign, but when i copied it, it looks like a multiplication dot thingy

OpenStudy (anonymous):

a = 7i - 4j, b = 4i + 3j he is asking for a multiplication then as we know that i and j are perpendicular to each other and their product becomes zero, ie i*j = 0 and i*i = 1 and j*j =1 as i and j are unit vectors now a*b = (7*4)(i*i) + (7*3)(i*j) + (-4*4)(j*i) + (-4*3)(j*j) a*b = 28 + 0 +0 =12 a*b = 28-12 a*b = 16

OpenStudy (anonymous):

so 16 would be the final answer??

OpenStudy (anonymous):

when we multiply ( scalar multiplication i mean just normal multiplication) two unit vectors then vector form vanishes yes, 16 is the final answer

OpenStudy (e.cociuba):

Finally! Whoo *dances in circle*. Ahem! Anyways, I'm sorry for the mistake, thanks @SandeepReddy ;)

OpenStudy (anonymous):

@ecociuba, \(a\cdot b\) is the dot product of \(a\) and \(b\). If \(a,b\in\mathbb{R}^n\) (meaning \(a=(a_1,a_2,...,a_n)\)), then \[a\cdot b=\sum_{k=1}^na_kb_k\]

OpenStudy (anonymous):

its ok, im just glad i got the answer right in the first place :p

OpenStudy (anonymous):

No problem e.cociuba :)

OpenStudy (anonymous):

thank you both so much! i wish i could give out 2 medals

OpenStudy (anonymous):

or we can just by heart form @SithsAndGiggles gave you. just multiplying corresponding numbers of unit vectors in this case i and j

OpenStudy (e.cociuba):

Thanks @SithsAndGiggles :) As you can see, my skills in pre-calc are a bit rusty. Anyways, sorry for the mistakes done and thanks for the corrections all :)